Ann lived in Bristol and Amy lived in Fairvale. At 2:00 P.M. they left their respective towns and walked in opposite directions, with Amy walking twice as fast an Ann. By 4:00 P.M., they were 66km apart. if,instead, they had walked from the towns toward each other, they would have been 12km apart at 3:00 P.M. How far apart are the towns?
Answer by mananth(16949)

Ann lived in Bristol and Amy lived in Fairvale. At 2:00 P.M. they left their respective towns and walked in opposite directions, with Amy walking twice as fast an Ann. By 4:00 P.M., they were 66km apart. if,instead, they had walked from the towns toward each other, they would have been 12km apart at 3:00 P.M. How far apart are the towns?
let the distance between towns be x miles
They move away from each other.
time = 2 hours
Distance = 66 km
Ann speed = y
Amy speed = 2y
Add up the speeds
=3y
66/3y=2
3y=66/2
3y=33
y=11 Ann speed
Amy speed = 22 kph
Their combined speed = 33 kph
They are 12 km away after 1 hour
so the distance between the towns = 33 +12 = 45 km